# Training-Free Label Space Alignment for Universal Domain Adaptation

### v1

@@ 3.2 Overview
$\bullet$ Step 1. The goal of this step is to filter the discovered labels by explicitly identifying and removing lexical synonyms of source labels using the WordNet database (Section. 3.3 and Fig. 4a).
$\bullet$ Step 2. The objective of this step is to remove discovered labels that are semantically ambiguous or resemble the source labels by identifying a similar label set in CLIP’s joint embedding space (Section 3.4 and Fig. 4b).
@@
$\bullet$ Step 3. To address rare instances missed in the previous phase, we removed labels based on their occurrence frequency, operating under the assumption that infrequent target private candidate labels were likely incorrect target private labels (Section 3.5 and Fig. 4c).
$\bullet$ Step 4. This step aims to fully adapt the zero-shot model to the target domain, thereby enhancing its performance beyond zero-shot inference. This enhancement is achieved by utilizing a universal classifier weighted by CLIP text embeddings (Section 3.6 and Fig. 4d).
@@ 3.3 Synonym Label Alignment
In this subsection, we introduce a method for generating raw labels and eliminating synonyms. To acquire the discovered labels, we employ a generative VLM to generate labels for unknown target images. Furthermore, we use WordNet (Miller et al., 1990) to remove semantically ambiguous labels (i.e., synonyms) from the discovered labels, ensuring their alignment with the source label space, as illustrated in Fig. 4(a).
Discovering labels on unlabeled target images. To ensure reliable labels for the input images, we query each target sample $x^{t}_{i}$ using the BLIP-VQA model (Li et al., 2022, 2023). Specifically, we utilize $k=5$ prompts that are semantically equivalent yet lexically diverse, as presented in Table  14. The design of these prompts is informed by the principles of prompt engineering in CLIP (Radford et al., 2021), employing multiple textual variants to reduce prompt sensitivity. Consider prompts like “What is the exact name of the object in the painting?” and “Identify the object in this photo with its exact name,” which share identical semantics but vary in phrasing. If the model’s response was overly specific (e.g., detailing a color) or did not correspond to a single valid label, we omitted it from subsequent majority voting. We determined the final discovered label by applying majority voting across $k$ responses. Although a comprehensive ablation study across different $k$ values was not performed, preliminary observations suggested that $k=5$ struck a balance between computational efficiency and stability; hence, we consistently employed this setting throughout our experiments.
Synonym label alignment using WordNet. WordNet (Miller et al., 1990) is a widely utilized lexical database that provides taxonomic relationships among words. Path-based similarity measures are computed by assessing the distances between two synsets within the hierarchy. Previous studies have applied WordNet to visual recognition (Ge et al., 2023) to explore parent–child relationships between labels. Additionally, WordNet can discern different lexical forms of a word, aiding in identifying synonymous relationships. Consequently, we employed WordNet as an essential filtering step to eliminate noisy labels in $R$.
@@ 3.4 Semantic Label Alignment
In this subsection, we detail the removal of hypernyms, hyponyms, incorrect (noisy) labels of the source class, and synonyms not present in the WordNet database. We employed semantic similarity to evaluate the relationship between the text representations of the source and candidate labels, based on their embedding distances from the input images. If candidate labels have a similar proximity to the image as the source labels, they are deemed identical to the corresponding source label, as displayed in Fig. 4b.
@@
where $N$ denotes the batch size. The left part of $S$ indicates the cosine similarity between the image embeddings and the source label embeddings, whereas the right part corresponds to the cosine similarity between the image embeddings and the filtered discovered label embeddings.
@@
Identification of Target Private Labels from Filtered Discovered Labels. Using semantic similarity, we aim to determine whether an input image belongs to the source or target private classes. An overview of this procedure is presented in Fig. 5. When determining whether a label originates from a source private or target, we found that naive alignment methods—such as top-1 prediction-based alignment and margin-based alignment—performed poorly in our experiments (see Table 8). For instance, when ‘backpack’ is the source label and ‘bag’ is the discovered label (Fig. 2c), ‘bag’ and ‘backpack’ are not synonyms in WordNet; however, within the Office31 context, they effectively represent the same class. Treating them as distinct categories leads to numerous ‘backpack’ samples being misclassified as ‘bag,’ resulting in artificial label splits. In such cases, top-1 predictions may alternate between ‘backpack’ and ‘bag’ due to uncontrollable image conditions, without signifying any meaningful semantic difference. These semantically ambiguous labels might be included among certain target private labels. To address this, the margin score between the top-1 and top-2 predictions can evaluate the distance between source labels and the filtered discovered labels. However, this method relies on a predefined fixed threshold to assess reliability, ignoring class similarities that may vary adaptively across datasets and domains.
@@
$\bullet$ Criterion 1. (Maximum-gap-based thresholding) A prominent gap in similarity scores was identified among the top $k$ rankings. The labels preceding this gap accurately correspond to multiple image predictions. Utilizing this criterion, we establish the meaningful prediction set $\mathcal{C}$. The threshold is defined as follows:
@@
$\bullet$ Criterion 2. (Average score-based thresholding) Similarly, we define the prediction set as comprising scores that exceed the average score of the top $k$ predictions. The subsequent equation specifies the threshold based on the average of the top $k$ scores.
@@
Empirically, we maximize the prediction set size by selecting the lower of the two criteria. This approach enhances the inclusion of source labels within the label set, thereby reducing the likelihood of ambiguous labels being identified as private target candidates. In other words, by selecting the smaller threshold between $\tau_{\text{gap}}$ and $\tau_{\text{avg}}$, the criterion for classifying a sample as target private becomes more stringent. The threshold $\tau_{\text{set}}$ and prediction set $\mathcal{C}$ for each instance are formulated as follows:
@@
where $\mathbf{t}$ represents the text comprising extensions of source labels and filtered discovered labels, and $I^{i}_{k}$ denotes the top-$k$ indices for the $i^{\text{th}}$ sample. Utilizing the selected prediction set $\mathcal{C}$, we perform a source label presence check to achieve semantic label alignment. Specifically, if a source label is present in $\mathcal{C}^{i}$ (Fig. 5a, sample $i$ is classified as a source sample. When it is absent from the prediction set $\mathcal{C}$, sample $i$ is identified as a target private sample (Fig. 5b).
@@ 3.5 Frequency-based Noisy Candidate Filtering
Noisy and incorrectly predicted private target candidate labels may still persist. To address this, we developed a method that filters these noisy labels by analyzing their frequencies, as shown in Fig. 4 (c). We hypothesized that low-frequency classes are likely noisy labels in the target domain. To precisely measure label frequency, we employed a frequency bank. For samples classified as source classes, the top-1 prediction $p_{i}$was stored in the frequency bank as a reference standard (Fig. 5a). Similarly, for samples classified as target private classes, the discovered label $r_{i}$is added to the frequency bank even if $r_{i}\notin\mathcal{C}^{i}$(Fig. 5b). The rationale for storing $r_{i}$, the label identified by BLIP, instead of $p_{i}$, the closest label in the CLIP space, is detailed in Appendix 11. The frequency bank is defined as $F=\{c:f_{c}\}$, where $f_{c}$ represents the frequency of class $c$ after semantic label alignment, including both source and filtered discovered labels. Considering the presence of noisy labels in the frequency bank, we used the count of the most frequent class rather than the median or mean of the label frequency distribution. We set a threshold $\tau_{\text{freq}}$as the fixed number of samples in the most frequent class. This scaling factor, denoted by $\epsilon$, is a dataset-agnostic parameter. Finally, the predicted target private labels $\mathcal{C}_{p}$ are derived from the raw discovered labels $R$ as follows:
@@
We employ an empirically determined value of $\epsilon=0.01$ across all datasets, as presented in Table 10.
By integrating complementary signals—lexical similarity validated with WordNet, semantic similarity calculated within a joint image–text embedding space, and a final statistical test that eliminates residual noisy discovered labels—our pipeline outperforms naïve filtering significantly, as demonstrated in Table 7.
@@ 3.6 Self-Training with Universal Classifier
In UniDA, the prior typically uses randomly initialized prototype-based classifiers (Li et al., 2021a; Chang et al., 2022; Lai et al., 2023). However, this method complicates determining the number of target clusters, which corresponds to the number of private target classes. In contrast, we utilize a universal classifier that automatically determines the number of target private classes, as depicted in Fig. 4d. Furthermore, we applied self-training, a prevalent technique in source-free domain adaptation Liang et al. (2020); Zhang et al. (2023); Karim et al. (2023); Tang et al. (2024). This method reduces bias toward the source domain, offering an advantage over approaches that depend on visual-space domain alignment and source-supervised learning. Additionally, we selected reliable pseudo-labels based on the top-$k$ confidence samples per class and employed a student-teacher framework with EMA updates.
@@
Reliable Pseudo Label Selection and Self-Training. To leverage pseudo-labels, we trained a student model to enhance the teacher model’s performance using delayed updates. However, naive pseudo-labeling tends to select only easily classifiable classes, leading to overfitting of these simpler categories. To address this, we introduce a balanced pseudo-label selection strategy that improves the generalization of the predictions of the pretrained model. Rather than requiring a large dataset, our approach prioritizes maximum diversity by selecting the top-k% most confident pseudo-labels per class, as proposed by Zara et al. (2023). Nonetheless, even with a top-k% threshold, more prevalent classes are disproportionately selected, which can degrade performance. To mitigate this, we define a balanced $k$ value to select only the top-$k$ most confident samples for each class. Specifically, $k$ is calculated as $k=\text{min}\{n_{c_{i}},N/n_{c}\}$, where $n_{c}$ is the number of unique predicted classes and $n_{c_{i}}$ is the number of predictions for class $c_{i}$.
To ensure robust self-training, we apply Exponential Moving Average (EMA) updates to the teacher model, defined by $\theta_{t}=(1-\alpha)\theta_{s}+\alpha\theta_{t}$. Here, $\theta_{t}$ and $\theta_{s}$ represent the parameters of the teacher and student models, respectively, effectively mitigating confirmation bias. The student model updates with each iteration, whereas the teacher model updates once per epoch, as defined by the preceding equation. This configuration is crucial for the performance of the model (Section 5). Additionally, we implemented adapter tuning to limit further tuning capabilities.
Consequently, we performed self-training using the following loss function: Let $M$ represent the set of samples employed as pseudo-labels, $\hat{y}^{T}_{t,i}$denotes the output of the teacher model, and $\hat{y}^{S}_{t,i}$represents the output of the student model.
@@
Table 1: Split-setting of our experiments. (Source Private / Shared)
@@
4 Experiments
Datasets
We utilize four benchmark datasets for visual recognition domain adaptation: Office31 (O), Office-Home (OH), VisDA (VD), and DomainNet (DN). Office31 (Saenko et al., 2010) is a small-scale benchmark dataset comprising images from 31 categories distributed across three domains: Amazon (2,817 images), Dslr (498 images), and Webcam (795 images). Office-Home (Venkateswara et al., 2017) is a benchmark dataset designed for domain adaptation, featuring images from 65 categories across four domains: Art (9,655 images), Clipart (9,516 images), Product (15,688 images), and Realworld (14,840 images). VisDA (Peng et al., 2017) is a benchmark dataset for domain adaptation, containing images from 12 labels distributed across two domains: Synthetic (152,397 images) and Real (55,088 images). DomainNet (Peng et al., 2019) is a large-scale benchmark dataset for domain adaptation, notable for its extensive size. It includes images from 345 labels distributed across six domains: Clipart, Infograph, Painting, QuickDraw, Real, and Sketch. The sheer scale of DomainNet makes it a particularly challenging and comprehensive dataset for evaluating domain adaptation methods. We conducted experiments using only three domains from DomainNet that are commonly employed in existing methods for UniDA - Painting, Sketch, and Real.
Baselines
@@ 4 Experiments
Implementation details
We use CLIP (Radford et al., 2021) backbone with ViT-B/16 (Dosovitskiy et al., 2020) as our image encoder. For the generative VLM, we utilize BLIP (Li et al., 2022), a model pretrained on VQAv2 (Goyal et al., 2017) within the LAVIS library [^1] https://github.com/salesforce/LAVIS. This model incorporates a visual encoder based on ViT-Base (Dosovitskiy et al., 2020) and employs a text encoder and decoder that adhere to the BERT (Devlin et al., 2018) architecture. Most of the baseline code and pipelines are based on the uniood repository [^2] https://github.com/szubing/uniood.
@@
In the self-training phase, we conducted experiments with a batch size of 64, a learning rate of 0.01, and an adapter with bottleneck dimensions of r=64, adhering to the default settings of AdaptFormer (Chen et al., 2022b). All baseline methods were implemented with identical iterations, batch sizes, and learning rates using a frozen backbone, as described in Deng and Jia (2023). The iterations were set to 2,500; 5,000; and 10,000, depending on the task setting, as presented in Table 2, following (Deng and Jia, 2023). Our methods required approximately 30 min per domain, allowing us to complete all experiments on the entire dataset within a day.
Table 2: Training iterations on different split settings. We follow the same setting in (Deng and Jia, 2023). Since we use batch size 64, which is twice the original value, we reduce the number of iterations accordingly by half.
@@
Evaluation metric
We use the existing two evaluation metrics for UniDA in addition to classification accuracy. H-score (Fu et al., 2020) is the harmonic mean of the instance accuracy on common class $a_{C}$ and accuracy on the single private class $a_{\bar{C_{t}}}$. This metric effectively captures the trade-off between known-class accuracy and out-of-class detection accuracy. Moreover, we also use the H3-score. The H3-score is first proposed by UniOT (Chang et al., 2022) and extends the H-Score by incorporating the Normalized Mutual Information (NMI), a well-known measure that indicates how well the target private clusters are formed. In the PDA and CDA settings where target private instances do not exist, the average class accuracy is used in place of the H-score. Similarly, the $H^{3}$ score is also substituted with this measure.

### v2

@@ 3.2 Overview
$\bullet$ Step 1. The goal of this step is to filter the discovered labels by explicitly identifying and removing lexical synonyms of source labels using the WordNet database (Sec. 3.3 and Fig. 4a).
$\bullet$ Step 2. The objective of this step is to remove discovered labels that are semantically ambiguous or resemble the source labels by identifying a similar label set in CLIP’s joint embedding space (Sec. 3.4 and Fig. 4b).
@@
$\bullet$ Step 3. To address rare instances missed in the previous phase, we removed labels based on their occurrence frequency, operating under the assumption that infrequent target private candidate labels were likely incorrect target private labels (Sec. 3.5 and Fig. 4c).
$\bullet$ Step 4. This step aims to fully adapt the zero-shot model to the target domain, thereby enhancing its performance beyond zero-shot inference. This enhancement is achieved by utilizing a universal classifier weighted by CLIP text embeddings (Sec. 3.6 and Fig. 4d).
@@ 3.3 Synonym Label Alignment
In this subsection, we introduce a method for generating raw labels and eliminating synonyms. To acquire the discovered labels, we employ a generative VLM to generate labels for unknown target images. Furthermore, we use WordNet (Miller et al., 1990) to remove semantically ambiguous labels (i.e., synonyms) from the discovered labels, ensuring their alignment with the source label space, as illustrated in Fig. 4a.
Discovering Labels on Unlabeled Target Images. To ensure reliable labels for the input images, we query each target sample $x^{t}_{i}$ using the BLIP-VQA model (Li et al., 2022, 2023). Specifically, we utilize $k=5$ prompts that are semantically equivalent yet lexically diverse, as presented in Tab. 14. The design of these prompts is inspired by the principles of prompt engineering in CLIP (Radford et al., 2021), employing multiple textual variants to reduce prompt sensitivity. Consider prompts like “What is the exact name of the object in the painting?” and “Identify the object in this photo with its exact name,” which share identical semantics but vary in phrasing. The final label was determined via majority voting over $k$ responses. Responses that are overly specific (e.g., specifying a color) or do not correspond to a single valid label are filtered out using a rule-based algorithm. Although a comprehensive sensitivity analysis across different $k$ values was not performed, preliminary observations suggested that $k=5$ struck a balance between computational efficiency and stability; hence, we consistently employed this setting throughout our experiments.
Synonym Label Alignment using WordNet. WordNet (Miller et al., 1990) is a widely utilized lexical database that provides taxonomic relationships among words. Path-based similarity measures are computed by assessing the distances between two synsets within the hierarchy. Consequently, we employed WordNet as an essential filtering step to eliminate noisy labels in $R$.
@@ 3.4 Semantic Label Alignment
This subsection elaborates on eliminating semantically ambiguous labels, such as hypernyms, hyponyms, and non-WordNet synonyms. We employed semantic similarity to evaluate the relationship between the text representations of the source and candidate labels, based on their embedding distances from the input images. If candidate labels have a similar proximity to the image as the source labels, they are deemed identical to the corresponding source label, as displayed in Fig. 4b.
@@
where $N$ denotes the batch size. The left portion of $\mathbf{S}$ reflects the cosine similarity between the embedding of the images and source labels, while the right portion represents the cosine similarity between images and the filtered discovered labels.
@@
Identification of Target Private Labels from Filtered Discovered Labels. Using semantic similarity, we aim to determine whether an input image belongs to the source or target private classes. An overview of this procedure is presented in Fig. 5. When determining whether a label originates from a source-private or target, we found that naive alignment methods—such as top-1 prediction-based alignment and margin-based alignment—performed poorly in our experiments (see Tab. 8). For instance, when ‘backpack’ is the source label and ‘bag’ is the discovered label (Fig. 2c), ‘bag’ and ‘backpack’ are not synonyms in WordNet; however, within the Office31 context, they effectively represent the same class. Treating them as distinct categories leads to numerous ‘backpack’ samples being misclassified as ‘bag’, resulting in artificial label splits. In such cases, top-1 predictions alternate between ‘backpack’ and ‘bag’ due to uncontrollable image conditions, without signifying any meaningful semantic difference. These semantically ambiguous labels might be included among certain target private labels. To address this, the margin value between the top-1 and top-2 predicted probabilities can evaluate the distance between source labels and the filtered discovered labels. However, this method relies on a predefined fixed threshold to assess reliability, ignoring class similarities that often vary adaptively across datasets and domains.
@@
$\bullet$ Criterion 1: Maximum-gap-based thresholding. A prominent gap in similarity scores was identified among the top $k$ rankings. The labels preceding this gap accurately correspond to multiple image predictions. Utilizing this criterion, we establish the meaningful prediction set $\mathcal{C}$. The threshold is defined as follows:
@@
$\bullet$ Criterion 2: Average score-based thresholding. We define the prediction set as comprising scores that exceed the average score of the top $k$ predictions. The threshold is defined as the average of the top $k$ scores:
@@
Empirically, we maximize the prediction set size by selecting the lower of the two criteria (i.e., $\tau^{i}_{gap}$ and $\tau^{i}_{avg}$). This approach enhances the inclusion of source labels within the label set, thereby reducing the likelihood of ambiguous labels being identified as private target candidates. In other words, by selecting the smaller threshold between $\tau_{\text{gap}}$ and $\tau_{\text{avg}}$, the criterion for classifying a sample as target private becomes more stringent. The threshold $\tau_{\text{set}}$ and prediction set $\mathcal{C}$ for each instance are formulated as follows:
@@
where $\mathbf{t}$ represents the text comprising extensions of source labels and filtered discovered labels, and $I^{i}_{k}$ denotes the top-$k$ indices for the $i^{\text{th}}$ sample. Utilizing the selected prediction set $\mathcal{C}$, we perform a source label presence check to achieve semantic label alignment. Specifically, if a source label is present in $\mathcal{C}^{i}$ (Fig. 5a), sample $i$ is classified as a source sample. When it is absent from the prediction set $\mathcal{C}$, sample $i$ is identified as a target private sample (Fig. 5b).
@@ 3.5 Frequency-based Noisy Candidate Filtering
Noisy and incorrectly predicted private target candidate labels may still persist. To address this, we developed a method that filters these noisy labels by analyzing their frequencies, as shown in Fig. 4c. We hypothesized that low-frequency classes are likely noisy labels in the target domain. To precisely measure label frequency, we employed a frequency bank. For samples classified as source classes, the top-1 prediction $p_{i}$ is stored in the frequency bank as a reference standard (Fig. 5a). Similarly, for samples classified as target private classes, the discovered label $r_{i}$ is added to the frequency bank even if $r_{i}\notin\mathcal{C}^{i}$(Fig. 5b). The rationale for storing $r_{i}$, the label identified by BLIP, instead of $p_{i}$, the closest label in the CLIP space, is detailed in Appendix D. The frequency bank is defined as $F=\{c:f_{c}\}$, where $f_{c}$ represents the frequency of class $c$ after semantic label alignment, including both source and filtered discovered labels. Considering the presence of noisy labels in the frequency bank, we used the count of the most frequent class rather than the median or mean of the label frequency distribution. We set a threshold $\tau_{\text{freq}}$ as the fixed number of samples in the most frequent class. This scaling factor, denoted by $\epsilon$, is a dataset-agnostic parameter. Finally, the predicted target private labels $\mathcal{C}_{p}$ are derived from the raw discovered labels $R$ as follows:
@@
We employ an empirically determined value of $\epsilon=0.01$ across all datasets, as presented in Tab. 10.
By integrating complementary components of the pipeline—lexical filtering validated through WordNet, semantic filtering within a joint image–text embedding space, and a frequency-based filtering step that removes residual noisy discovered labels depending on dataset scale—our method achieves strong complementarity among the three, as evidenced in Tab. 7.
@@ 3.6 Self-Training with Universal Classifier
In UniDA, the prior typically uses randomly initialized prototype-based classifiers (Li et al., 2021a; Chang et al., 2022; Lai et al., 2023). However, this method complicates determining the number of target clusters, which corresponds to the number of private target classes. In contrast, we utilize a universal classifier that automatically determines the number of target private classes, as depicted in Fig. 4d. Furthermore, we applied self-training, a prevalent technique in source-free domain adaptation (Liang et al., 2020; Zhang et al., 2023; Karim et al., 2023; Tang et al., 2024). This method reduces bias toward the source domain, offering an advantage over approaches that depend on visual-space domain alignment and source-supervised learning. Additionally, we selected reliable pseudo labels based on the top-$k$ confidence samples per class and employed a student-teacher framework with EMA updates.
@@
Reliable Pseudo Label Selection. To leverage pseudo-labels, we trained a student model to enhance the teacher model’s performance using delayed updates. However, naive pseudo-labeling tends to select only easily classifiable classes, leading to overfitting of these simpler categories. To address this, we introduce a balanced pseudo-label selection strategy that improves the generalization of the predictions of the pretrained model. Rather than requiring a large dataset, our approach prioritizes maximum diversity by selecting the top-k (%) most confident pseudo-labels per class, as proposed by Zara et al. (2023). Nonetheless, even with a top-k (%) threshold, more prevalent classes are disproportionately selected, which can degrade performance. To mitigate this, we define a balanced $k$ value to select only the top-$k$ most confident samples for each class. Specifically, $k$ is calculated as $k=\text{min}\{n_{c_{i}},N/n_{c}\}$, where $n_{c}$ is the number of unique predicted classes and $n_{c_{i}}$ is the number of predictions for class $c_{i}$.
Robust Self-Training. To ensure robust self-training, we apply Exponential Moving Average (EMA) updates to the teacher model, defined by $\theta_{t}=(1-\alpha)\theta_{s}+\alpha\theta_{t}$. Here, $\theta_{t}$ and $\theta_{s}$ represent the parameters of the teacher and student models, respectively, effectively mitigating confirmation bias. The student model updates with each iteration, whereas the teacher model updates once per epoch, as defined by the preceding equation. This configuration is crucial for the performance of the model (Sec. 5). Additionally, we implemented adapter tuning to limit further tuning capabilities. We then perform self-training using the following loss function: Let $M$ represent the set of samples employed as pseudo-labels, $\hat{y}^{T}_{t,i}$denotes the output of the teacher model, and $\hat{y}^{S}_{t,i}$represents the output of the student model.
@@
4 Experiments
Datasets. We utilize four benchmark datasets for visual recognition domain adaptation: Office31 (O), Office-Home (OH), VisDA (VD), and DomainNet (DN). Office31 (Saenko et al., 2010) is a small-scale benchmark dataset comprising images from 31 categories distributed across three domains: Amazon (2,817 images), Dslr (498 images), and Webcam (795 images). Office-Home (Venkateswara et al., 2017) is a benchmark dataset designed for domain adaptation, featuring images from 65 categories across four domains: Art (9,655 images), Clipart (9,516 images), Product (15,688 images), and Realworld (14,840 images). VisDA (Peng et al., 2017) is a benchmark dataset for domain adaptation, containing images from 12 labels distributed across two domains: Synthetic (152,397 images) and Real (55,088 images). DomainNet (Peng et al., 2019) is a large-scale benchmark dataset for domain adaptation, notable for its extensive size. It includes images from 345 labels distributed across six domains: Clipart, Infograph, Painting, QuickDraw, Real, and Sketch. The sheer scale of DomainNet makes it a particularly challenging and comprehensive dataset for evaluating domain adaptation methods. We conducted experiments using only three domains from DomainNet that are commonly employed in existing methods for UniDA - Painting, Sketch, and Real.
Table 1: Split-setting of our experiments. Details of experimental split settings indicating the counts of source-private and shared classes (Source-Private / Shared) in each dataset under different adaptation scenarios (open-partial, open, closed, partial).
@@ 4 Experiments
Implementation Details. We use CLIP (Radford et al., 2021) backbone with ViT-B/16 (Dosovitskiy et al., 2020) as our image encoder. For the generative VLM, we utilize BLIP (Li et al., 2022), a model pretrained on VQAv2 (Goyal et al., 2017) within the LAVIS library [^1] https://github.com/salesforce/LAVIS. This model incorporates a visual encoder based on ViT-Base (Dosovitskiy et al., 2020) and employs a text encoder and decoder that adhere to the BERT (Devlin et al., 2018) architecture. Most of the baseline code and pipelines are based on the uniood repository [^2] https://github.com/szubing/uniood.
Our experimental class split settings follow the standard protocol used by many existing UniDA approaches (Saito et al., 2020; Saito and Saenko, 2021; Chang et al., 2022; Lu et al., 2024; Qu et al., 2024b, c; Deng and Jia, 2023), as detailed in Tab. 1.
@@
In the self-training phase, we conducted experiments with a batch size of 64, a learning rate of 0.01, and an adapter with bottleneck dimensions of r=64, adhering to the default settings of AdaptFormer (Chen et al., 2022b). All baseline methods were implemented with identical iterations, batch sizes, and learning rates using a frozen backbone, as described in Deng and Jia (2023). The iterations were set to 2,500; 5,000; and 10,000, depending on the task setting, as presented in Tab. 2, following (Deng and Jia, 2023). Our methods required approximately 30 minutes per domain, allowing us to complete all experiments on the entire dataset within a day.
Table 2: Training iterations on different split settings. We follow the same setting in (Deng and Jia, 2023). Since we use a batch size 64, which is twice the original value, we reduce the number of iterations accordingly by half.
@@
Evaluation Metric. We use the existing two evaluation metrics for UniDA in addition to classification accuracy. H-score (Fu et al., 2020) is the harmonic mean of the instance accuracy on common class $a_{C}$ and accuracy on the single private class $a_{\bar{C_{t}}}$. This metric effectively captures the trade-off between known-class accuracy and out-of-class detection accuracy. Moreover, we also use the H3-score. The H3-score is first proposed by UniOT (Chang et al., 2022) and extends the H-Score by incorporating the Normalized Mutual Information (NMI), a well-known measure that indicates how well the target-private clusters are formed. In the PDA and CDA settings where target-private instances do not exist, the average class accuracy is used in place of the H-score. Similarly, the $H^{3}$ score is also substituted with this measure.